Order: Chloramphenicol (Chloromycetin) 12.5mg/kg/q6h IV.

Infant weighs 3000 grams.

Available: Chloramphenicol 100mg/mL

How many mL's will the nurse administer for each dose?

Round your answer to the nearest 10th.

________ mLs

pls show all work, tu
Answer #1

Prescribed a dose of chloramphenicol 12.5mg/kg 6th hourly through intravenous

Infant weight is 3000 gms = 3 kgs (divide grams with 1000 to get kg value)

The dose is 3 \times 12.5mg = 37.5 mg

The required volume = required dose \times available volume \div available dose

                                  = 37.5 \times 1 ml \div 100 mg

                                  = 0.37 ml (can be rounded to 0.4)

The nurse should administer 0.4 ml of chloramphenicol for each dose.